# MEDIC: a network for monitoring data quality in collider experiments

### v1

@@ 5 Discussion and Conclusion
The advantage of this simulation-driven approach lies in its flexibility. One can simulate a variety of detector conditions, which allows ML models to learn robust representations of anomalous behaviors. In the present study, MEDIC is trained on particle-level kinematics obtained from Delphes simulations [deFavereau:2013fsa]. Even though only large-scale detector components were switched off in the scenarios presented here, the data was generated such that any temporal window could contain any combination of the four simulated situations. Notably, MEDIC was not explicitly trained for binary classification between normal and anomalous runs, nonetheless, it demonstrates strong predictive performance for that as well.
@@
It should be emphasized that Delphes provides only particle-level outputs, not detailed electronic-level signals. Its digitization process is inherently limited given its modular and fast-simulation framework that relies on parameterization of the overall detector response instead of simulating the underlying physics processes in detail. This simplification limits the resolution and number of features available to the network, which occasionally leads to ambiguities between normal and anomalous events as can be observed from the confusion matrices in figure 5. A more detailed simulation using Geant4 [GEANT4:2002zbu], would likely improve the model accuracy as Geant4 performs a detailed simulation with precise position and timing information of energy deposits. It also converts raw energy deposits into realistic detector signals, accounting for specific hardware effects coming from signal generation, transport and electronic response combined with detector resolution effects. This translates into more information rich features using Geant4 that might reflect in improved accuracy. Despite this limitations, our results indicate that a ML model trained solely on Delphes-level data can still successfully detect anomalies and even identify their sources.
@@
This work demonstrates the potential of simulation-driven DQM as a new paradigm for collider experiments. Any experiment can benefit from generating simulated reference runs and glitch scenarios prior to data-taking runs, enabling the development and validation of automated monitoring tools in advance. Although the current result is promising and a more detailed simulation with electronic input and perhaps with more data points might increase the accuracies of prediction, our approach is still intended as a human-in-the-loop system, where shifters are assisted by the network’s predictions rather than fully replaced. Such a pipeline allows for rapid anomaly detection and classification, providing actionable insights in real time and significantly aiding detector operations. This would be particularly valuable in the context of upcoming HL-LHC runs with their extremely high data volumes, where automated tools such as MEDIC could significantly enhance the efficiency of the shifter by rapidly flagging anomalies along with pointing out the source of error, thereby significantly reducing the manual load of real-time DQM.
@@ Acknowledgment
This work is supported by US NSF Award 2334265. We would like to thank Roy Cruz (UPR Mayagüez), Tetiana Mazurets (UPR Mayagüez) and Gabriele Benelli (Brown University) for illuminating discussions.
@@ Appendix A Dataset Generation
The first step in the data generation pipeline for MEDIC consisted of using MadGraph5_aMC@NLO [Alwall:2014hca] to generate $100,000$ proton-proton collision events to produce two jets with $\sqrt{s}=13$ TeV. The resulting file was then passed through Pythia8 [Bierlich:2022pfr] for parton showering and hadronization. Then Delphes [deFavereau:2013fsa, arghya_delphes] is used to simulate the 4 distinct CMS detector responses as mentioned at the end of section˜2 (corresponding $\eta$ distributions are shown in fig.˜2).

### v2

@@ 5 Discussion and Conclusion
The advantage of this simulation-driven approach lies in its flexibility. One can simulate a variety of detector conditions, which allows ML models to learn robust representations of anomalous behaviors. In the present study, MEDIC is trained on particle-level kinematics obtained from Delphes simulations [22]. Even though only large-scale detector components were switched off in the scenarios presented here, the data was generated such that any temporal window could contain any combination of the four simulated situations. Notably, MEDIC was not explicitly trained for binary classification between normal and anomalous runs, nonetheless, it demonstrates strong predictive performance for that as well.
@@
It should be emphasized that Delphes provides only particle-level outputs, not detailed electronic-level signals. Its digitization process is inherently limited given its modular and fast-simulation framework that relies on parameterization of the overall detector response instead of simulating the underlying physics processes in detail. This simplification limits the resolution and number of features available to the network, which occasionally leads to ambiguities between normal and anomalous events as can be observed from the confusion matrices in figure 5. A more detailed simulation using Geant4 [7], would likely improve the model accuracy as Geant4 performs a detailed simulation with precise position and timing information of energy deposits. It also converts raw energy deposits into realistic detector signals, accounting for specific hardware effects coming from signal generation, transport and electronic response combined with detector resolution effects. This translates into more information rich features using Geant4 that might reflect in improved accuracy. Despite this limitations, our results indicate that a ML model trained solely on Delphes-level data can still successfully detect anomalies and even identify their sources.
@@
This work demonstrates the potential of simulation-driven DQM as a new paradigm for collider experiments. Any experiment can benefit from generating simulated reference runs and glitch scenarios prior to data-taking runs, enabling the development and validation of automated monitoring tools in advance. Although the current result is promising and a more detailed simulation with electronic input and perhaps with more data points might increase the accuracies of prediction, our approach is still intended as a human-in-the-loop system, where shifters are assisted by the network’s predictions rather than fully replaced. In this framework, alerts are issued only after persistent anomalous behavior is observed over multiple consecutive windows, ensuring that human attention is drawn to sustained issues rather than simple statistical fluctuations. Such a pipeline allows for rapid anomaly detection and classification, providing actionable insights in real time and significantly aiding detector operations. This would be particularly valuable in the context of upcoming HL-LHC runs with their extremely high data volumes, where automated tools such as MEDIC could significantly enhance the efficiency of the shifter by rapidly flagging anomalies along with pointing out the source of error, thereby significantly reducing the manual load of real-time DQM.
While the present study focuses on a limited number of labeled detector configurations as a proof of concept, the MEDIC framework is intended to be extended to a wider range of detector conditions, including partially labeled or unlabeled scenarios. In a realistic DQM deployment, this does not require a single classifier with a very large number of output classes. Instead, the framework can scale naturally through hierarchical or modular strategies, where deviations from nominal detector behavior are detected first and then classified at the level of subsystems or detector regions. A key feature of this approach is that the training data are generated entirely through simulation by explicitly introducing controlled detector fault scenarios, rather than relying only on statistical fluctuations in real data, which distinguishes MEDIC from many traditional DQM methods.
In this spirit, MEDIC can be viewed as a modular architecture in which shared event-encoding layers are reused, while only the final classification head is adapted or replaced as new detector scenarios are introduced. This allows multiple binary or few-class classifiers, guided by historical knowledge of common failure modes, to operate in parallel and be extended incrementally without retraining the full network. The same framework can also be adapted to unsupervised or semi-supervised settings by learning a representation of nominal detector behavior and flagging persistent deviations in the latent space, enabling sensitivity to unforeseen failure modes. As with any simulation-driven approach, network performance may be affected by inaccuracies in detector modeling; a systematic study of such effects using dedicated Monte Carlo variations is therefore an important direction for future work and beyond the scope of the present study. The implementation associated with this work is publicly available in the MEDIC GitHub repository [13] and is provided to support transparency and reproducibility of the results presented here. The repository currently reflects an active research codebase, and future methodological developments discussed in this work will be incorporated within the same framework.
@@ Acknowledgment
This work is supported by U.S.A. National Science Foundation Award OAC-2334265. We would like to thank Roy Cruz (UPR Mayagüez), Tetiana Mazurets (UPR Mayagüez) and Gabriele Benelli (Brown University) for illuminating discussions.
@@ Appendix A Dataset Generation
The first step in the data generation pipeline for MEDIC consisted of using MadGraph5_aMC@NLO [9] to generate $100,000$ proton-proton collision events to produce two jets with $\sqrt{s}=13$ TeV. The resulting file was then passed through Pythia8 [16] for parton showering and hadronization. Then Delphes [22, 14] is used to simulate the 4 distinct CMS detector responses as mentioned at the end of section˜2 (corresponding $\eta$ distributions are shown in fig.˜2).